# Sonja Hohloch  &\&&  Pedro Santos

### v1

@@ 2.6. Almost toric manifolds
So called almost toric manifolds, as introduced by Symington [symington2002four], are a generalization of closed toric manifolds and will be defined in more detail below.
@@
From a toric base $(B,\mathcal{A},\mathcal{S})$ it is possible to reconstruct a toric fibration and a symplectic manifold by starting with a regular Lagrangian fibration over the base $(B,\mathcal{A})$ and then collapse certain fibers to lower dimensional tori so as to obtain the stratification $\mathcal{S}$. The collapsing of the fibers is achieved via boundary reduction, see Symington [symington2002four, Proposition $3.8$] for more details.
@@
Let $\pi:(M,\omega)\rightarrow(B,\mathcal{A},\mathcal{S})$ be a toric fibration and $\gamma$ the image in $B$ of a compact embedded curve with one endpoint $b_{1}$ in the $(n-1)$-dimensional stratum of $\partial_{R}B$ and such that $\gamma\setminus\{b_{1}\}\subset B_{r}$. Let $b_{2}\in B$ be the other endpoint of $\gamma$. The collapsing class with respect to $\gamma$ for the smooth component of $\partial_{R}B$ containing $b_{1}$ is defined as the class $a\in H_{1}(F_{b_{2}},\mathbb{Z})$ that spans the kernel of $i_{*}:H_{1}(F_{b_{2}},\mathbb{Z})\rightarrow H_{1}(\pi^{-1}(\gamma),\mathbb{Z})$ where $i$ is the inclusion map $i:F_{b_{2}}\rightarrow\pi^{-1}(\gamma)$. Accordingly, the collapsing covector with respect to $\gamma$ is the covector $v^{*}\in T^{*}_{b_{2}}B$ that determines vectors $v(x)\in T^{vert}_{x}M:=\ker(d_{x}\pi)$ for each $x\in\pi^{-1}(B_{r})$ such that the flow lines of this vector field represent $a$. (Notice $a$ and $v^{*}$ are well-defined up to sign and have coefficients in $\mathbb{Z}$.)
@@
(Symington, [symington2002four, Lemma 4.10]) Using the notation from Definition 2.9 suppose $\gamma:S^{1}\rightarrow B$ is a positively oriented loop based at $b\in B$ that is the boundary of a closed neighborhood of $s\in B$ containing $\eta\subset B$. Then the vanishing class, up to rescaling, is the unique class that is preserved by the monodromy along $\gamma$.
Let $\pi:(U,\omega)\rightarrow B$ be a Lagrangian fibered neighborhood of a nodal fiber in a $4$-manifold where $B$ is connected and (the value of) the node is $s\in B$. Let $\mathcal{A}$ be the affine structure on $B\setminus\{s\}$. The affine monodromy is non-trivial, due to the presence of the node (Symington, [symington2002four] Section $4.3$). Hence there is no immersion of $(B\setminus\{s\},\mathcal{A})$ into $(\mathbb{R}^{2},\mathcal{A}_{0})$. However, there is an affine immersion if one removes a curve from $B^{2}$ that connects the (value of the) node to the boundary since the remaining set is simply connected. We define the base diagram for the neighborhood of a node, as the image of an affine immersion of the complement of such a curve.
@@ 2.7. Nodal trades and slides
(Symington, [symington2002four, Proposition 6.2]) If two bases are related by a nodal slide then they define the same manifold equipped with isotopic symplectic structures. Furthermore, the manifolds are fiberwise symplectomorphic on the complement of a compact set belonging to the preimage of the eigenline.
@@
(Symington, [symington2002four, Lemma 6.3]) Let $(B,\mathcal{A},\mathcal{S})$ be an almost toric base and $R\subset B$ be the image of an embedded eigenray that connects a node $s\in B$ with a point $b$ in an edge $E\subset\partial_{R}B$ such that there are no other nodes on $R$. Let $v^{*}$, $w^{*}\in T_{b}(B)^{*}$ be the vanishing and collapsing covectors. If $v^{*}$ and $w^{*}$ span $\Lambda_{b}^{*}$ then there is an almost toric base $(B,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$ such that:
@@
(Symington [symington2002four, Theorem 6.5]) Suppose two almost toric bases that are related by a nodal trade. Then the symplectic manifolds induced by the almost toric bases are symplectomorphic. In fact, their symplectic structures are isotopic.
@@ 2.8. Semitoric systems
If $M$ is $4$-dimensional, semitoric systems were classified, first under some conditions, by Pelayo &\&& Vũ Ngǫc [pelayo2009semitoric, pelayo2011constructing] and then in full generality by Palmer & Pelayo & Xiudi [palmer2019semitoric]. The distinguishing property of a semitoric system is that all but maybe one of the integrals generate periodic flows. So toric systems are a special case of semitoric ones. We will consider examples in Sections 4.1 & 4.2 & 4.3.
A simple semitoric system is a semitoric system where each fiber contains at most one focus-focus singularity. Such systems are determined up to isomorphism, according to Pelayo &\&& Vũ Ngǫc [pelayo2009semitoric], by the following $5$ symplectic invariants:
@@
[…]
@@
(Vũ Ngǫc, [vu2007moment, Theorem 3.8]) Using the notation from above for a semitoric system $(M,\omega,(f_{1},f_{2}))$, for all $\vec{\epsilon}\in\{-1,+1\}^{m_{f}}$, there exists a homeomorphism $f_{\vec{\epsilon}}=(f_{\vec{\epsilon}}^{(1)},f_{\vec{\epsilon}}^{(2)})$ from $B$ to $f_{\vec{\epsilon}}(B)\subseteq{\mathbb{R}}^{2}$ such that
@@
Describing all possible choices (like for instance the signs $\epsilon_{j}$) by means of a group action allows to write the polytope invariant as equivalence class or orbit of this group action. Since we work later always with representatives, we refer for details to Vũ Ngǫc [vu2007moment].
@@ 2.9. Semitoric systems and almost toric fibrations
Let $(M,\omega,F)$ be a semitoric system. In particular, it induces an almost toric fibration and an almost toric base. Let $\Delta_{\vec{\epsilon}}$ be a representative of the polytope invariant for this semitoric system. Then, by Vũ Ngǫc [vu2007moment], the eigenline of a nodal value in $\Delta_{\vec{\epsilon}}$ is the vertical line that passes through the node. The affine monodromy can be represented, with respect to an integral affine basis of $\Delta_{\vec{\epsilon}}$, as
@@ 2.10. Symplectic reduction
In this section we recall the following symplectic reduction result from Abreu & Macarini [abreu2013remarks]: let $x\in\text{int}(P)$ and let $T_{x}:=\mu^{-1}(x)$ denote the corresponding $\mathbb{T}^{N}$ orbit (which is a Lagrangian torus in $M$). Let $K\subset\mathbb{T}^{N}$ be a subtorus of dimension $N-n\geq 0$ determined by an inclusion of Lie algebras $\imath:\mathbb{R}^{N-n}\rightarrow\mathbb{R}^{N}$. The momentum map for the induced action of $K$ on $M$ is given by
@@ 2.11. Displacing toric fibers
In McDuff [mcduff2011displacing] conditions are given to displace Lagrangian toric fibers of toric manifolds. In this section, we summarize the necessary notions and results of McDuff [mcduff2011displacing] which we will need later on.
The affine distance $d_{\text{aff}}(x,y)$ between two points $x$, $y$ on a line $L$ with rational slope is defined as the quotient of the Euclidean distance $d_{E}(x,y)$ and the minimum Euclidean distance from 00 to an integral point $p\in\mathbb{Z}^{n}$ on the line through 00 parallel to $L$. Equivalently, if $\phi$ is any integral affine transformation of $\mathbb{R}^{n}$ that takes $x,y$ to the first coordinate axis (the $x_{1}$-axis) then $d_{\text{aff}}(x,y)=d_{E}(\phi(x),\phi(y))$. An affine line $L=z+\mathbb{R}v$ in $\mathbb{R}^{n}$ is rational if the direction vector $v$ can be taken to be a primitive integral vector in $\mathbb{Z}^{n}$. Given an affine rational line $L$ with primitive direction $v\in\mathbb{Z}^{n}$, the affine distance $d_{\text{aff}}$ between two points $x,y\in L$ is $t\in\mathbb{R}$ such that $x-y=tv\in\mathbb{R}^{n}$. An affine hyperplane $A$ is called rational if it has a primitive integral normal vector $\eta$, i.e., if it is given by an equation of the form $\langle x,\eta\rangle=k$ where $k\in\mathbb{R}$ and $\eta$ is primitive and integral. The affine distance $d_{\lambda}(x,A)$ from a point $x$ to a rational affine hyperplane in the rational direction $\lambda$ is defined as
@@
Let $w$ be a point on a facet $F$, i.e., on a $(n-1)$-dimensional face of a rational polytope $\Delta$ and let $\lambda\in\mathbb{Z}^{n}$ be integrally transverse to $F$. The probe $p_{F,\lambda}(w)=:p_{\lambda}(w)$ with direction $\lambda\in\mathbb{Z}^{n}$ and initial point $w\in F$ is the half open line segment consisting of $w$ together with the points in $\text{int}(\Delta)$ that lie on the ray from $w$ in direction $\lambda$.
@@
(McDuff [mcduff2011displacing, Lemma 2.4]) Let $\Delta$ be a smooth momentum polytope associated with a symplectic toric manifold $(M,\omega,F)$, i.e., $\Delta=F(M)$. Let $u\in\text{int}(\Delta)$ lie on the probe $p_{F,\lambda}(w)$. If $w$ lies in the interior of its corresponding facet and if $u$ is less than halfway measured from the boundary along $p_{\lambda}(w)$, then the fiber $F^{-1}(u)$ is displaceable.
@@ 2.12. Special Hamiltonian diffeomorphisms of the disk
In this subsection, we recall a result from Abreu &\&& Borman &\&& McDuff [abreu2014displacing] that concerns Hamiltonian diffeomorphisms of the disk which we will need to displace focus-focus fibers in Section 3.3.
@@
For our purposes, we pick $a(s):=B-s$. For $r>0$, let $\mathbb{D}(r)\subset\mathbb{R}^{2}$ denote the $2$-dimensional disk of radius $r$ centered at the origin.
Lemma 2.24 (Abreu &\&& Borman &\&& McDuff, [abreu2014displacing, Lemma 6.4.2]).
@@
where $\mathbb{A}(b,c):=\mathbb{D}(b)\backslash\text{Int}(\mathbb{D}(b))$.
@@
If $A>\frac{B}{2}$, every circle of radius less than $B-A$ is displaced by $\rho$.
@@ 2.13. Symplectic quasi-states
Symplectic quasi-states are a useful tool to find nondisplaceable subsets in a symplectic manifold, and so are the notions of superheavy, heavy and pseudoheavy sets. Entov &\&& Polterovich [entov2009rigid, polterovich2006quasi] introduced the notions of heaviness and superheaviness. Furthermore, Kawasaki & Orita [kawasaki2021existence] added the notion of pseudoheaviness. The definition of a symplectic quasi-state and pseudoheavy, heavy, superheavy subsets is based on spectral invariants in Hamiltonian Floer theory (see Oh [oh2005construction]). We now recall in more detail some definitions and results from these papers that we need later on.
@@
[…]
@@
$|\zeta(H+K)-\zeta(H)-\zeta(K)|\leq C\sqrt{||\{H,K\}||}\quad\forall\ H,K\in C^{\infty}(M,\omega).$
@@
[…]
@@
[…]
@@
(Kawasaki &\&& Orita [kawasaki2021existence, Proposition 1.8], Entov &\&& Polterovich [entov2009rigid, Theorem $1.4$]) Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$:
@@
(Kawasaki &\&& Orita, [kawasaki2021existence, Theorem 1.7]) Let $(M,\omega,F)$ be an integrable system. Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$. Then there exists $y_{0}\in F(M)$ such that $F^{-1}(y_{0})$ is $\zeta$-pseudoheavy.
@@
Entov & Polterovich [polterovich2006quasi] showed the existence of a partial symplectic quasi-state for $2n$-dimensional symplectic manifolds which are rational and strongly semi-positive. Due to the work of Oh [oh2009floer] and Usher [dualityusher, spectralusher] the conditions can be dropped, see Entov [entovclarification] for more details. Note that all symplectic 4-manifolds are strongly semi-positive.
@@
(Entov &\&& Polterovich, [polterovich2006quasi, Theorem 2.1], Usher [dualityusher, spectralusher], Oh [oh2009floer]) Any integrable system $(M,\omega,F)$ on a symplectic $4$-manifold $(M,\omega)$ has a nondisplaceable fiber.
@@
(Polterovich &\&& Rosen, [polterovich2014function, Corollary 6.1.6]) Any stem is nondisplaceable by the group of symplectomorphisms of $(M,\omega)$.
Borman [borman2013quasi] showed that symplectic quasi-states behave well under symplectic reduction on a superheavy subset: let $(W,\omega)$ be a $2n$-dimensional closed symplectic manifold equipped with a smooth map $\Phi=(\Phi_{1},\cdot\cdot\cdot,\Phi_{k}):W\rightarrow\mathbb{R}^{k}$ and a regular level set $Z:=\Phi^{-1}(0)$. Suppose that all $\Phi_{i}$ mutually Poisson commute at each point in $Z$ and that $\Phi$ induces a free Hamiltonian $\mathbb{T}^{k}$-action on $Z$. Let $(M^{\text{red},0}:=Z/\mathbb{T}^{k},\omega^{\text{red},0})$ be the reduced space when performing symplectic reduction at 00. Denote by $\pi:Z\rightarrow M^{\text{red},0}$ the quotient map.
@@
(Borman, [borman2013quasi, Theorem 1.1]) Let $(W,\omega)$ be a $2n$-dimensional closed symplectic manifold and $Z$ has above. If $\zeta:C^{\infty}(W,\omega)\rightarrow\mathbb{R}$ is a symplectic quasi-state satisfying the PB-inequality and $Z$ is superheavy for $\zeta$, then $\zeta$ induces a symplectic quasi-state

### v2

@@ 2.6. Almost toric manifolds
So called almost toric manifolds, as introduced by Symington [Sym03], are a generalization of closed toric manifolds and will be defined in more detail below.
@@
From a toric base $(B,\mathcal{A},\mathcal{S})$ it is possible to reconstruct a toric fibration and a symplectic manifold by starting with a regular Lagrangian fibration over the base $(B,\mathcal{A})$ and then collapse certain fibers to lower dimensional tori so as to obtain the stratification $\mathcal{S}$. The collapsing of the fibers is achieved via boundary reduction, see Symington [Sym03, Proposition $3.8$] for more details.
@@
Let $\pi:(M,\omega)\rightarrow(B,\mathcal{A},\mathcal{S})$ be a toric fibration and $\gamma$ the image in $B$ of a compact embedded curve with one endpoint $b_{1}$ in the (n1)(n-1)-dimensional stratum of $\partial_{R}B$ and such that $\gamma\setminus\{b_{1}\}\subset B_{r}$. Let $b_{2}\in B$ be the other endpoint of $\gamma$. The collapsing class with respect to $\gamma$ for the smooth component of $\partial_{R}B$ containing $b_{1}$ is defined as the class $a\in H_{1}(F_{b_{2}},\mathbb{Z})$ that spans the kernel of $i_{*}:H_{1}(F_{b_{2}},\mathbb{Z})\rightarrow H_{1}(\pi^{-1}(\gamma),\mathbb{Z})$ where $i$ is the inclusion map $i:F_{b_{2}}\rightarrow\pi^{-1}(\gamma)$. Accordingly, the collapsing covector with respect to $\gamma$ is the covector $v^{*}\in T^{*}_{b_{2}}B$ that determines vectors $v(x)\in T^{vert}_{x}M:=\ker(d_{x}\pi)$ for each $x\in\pi^{-1}(B_{r})$ such that the flow lines of this vector field represent $a$. (Notice $a$ and $v^{*}$ are well-defined up to sign and have coefficients in $\mathbb{Z}$.)
@@
(Symington, [Sym03, Lemma 4.10]) Using the notation from Definition 2.9 suppose $\gamma:\mathbb{S}^{1}\rightarrow B$ is a positively oriented loop based at $b\in B$ that is the boundary of a closed neighborhood of $s\in B$ containing $\eta\subset B$. Then the vanishing class, up to rescaling, is the unique class that is preserved by the monodromy along $\gamma$.
Let $\pi:(U,\omega)\rightarrow B$ be a Lagrangian fibered neighborhood of a nodal fiber in a $4$-manifold where $B$ is connected and (the value of) the node is $s\in B$. Let $\mathcal{A}$ be the affine structure on $B\setminus\{s\}$. The affine monodromy is non-trivial, due to the presence of the node (Symington, [Sym03] Section $4.3$). Hence there is no immersion of $(B\setminus\{s\},\mathcal{A})$ into $(\mathbb{R}^{2},\mathcal{A}_{0})$. However, there is an affine immersion if one removes a curve from $B^{2}$ that connects the (value of the) node to the boundary since the remaining set is simply connected. We define the base diagram for the neighborhood of a node, as the image of an affine immersion of the complement of such a curve.
@@ 2.7. Nodal trades and slides
(Symington, [Sym03, Proposition 6.2]) If two bases are related by a nodal slide then they define the same manifold equipped with isotopic symplectic structures. Furthermore, the manifolds are fiberwise symplectomorphic on the complement of a compact set belonging to the preimage of the eigenline.
@@
(Symington, [Sym03, Lemma 6.3]) Let $(B,\mathcal{A},\mathcal{S})$ be an almost toric base and $R\subset B$ be the image of an embedded eigenray that connects a node $s\in B$ with a point $b$ in an edge $E\subset\partial_{R}B$ such that there are no other nodes on $R$. Let $v^{*}$, $w^{*}\in T_{b}(B)^{*}$ be the vanishing and collapsing covectors. If $v^{*}$ and $w^{*}$ span $\Lambda_{b}^{*}$ then there is an almost toric base $(B,\mathcal{A}^{\prime},\mathcal{S}^{\prime})$ such that:
@@
(Symington [Sym03, Theorem 6.5]) Suppose two almost toric bases that are related by a nodal trade. Then the symplectic manifolds induced by the almost toric bases are symplectomorphic. In fact, their symplectic structures are isotopic.
@@ 2.8. Semitoric systems
If $M$ is $4$-dimensional, semitoric systems were classified, first under some conditions, by Pelayo &\& Vũ Ngǫc [PV09, PV11] and then in full generality by Palmer & Pelayo & Xiudi [PPT24]. The distinguishing property of a semitoric system is that all but maybe one of the integrals generate periodic flows. So toric systems are a special case of semitoric ones. We will consider examples in Sections 4.1 & 4.2 & 4.3.
A simple semitoric system is a semitoric system where each fiber contains at most one focus-focus singularity. Such systems are determined up to isomorphism, according to Pelayo &\& Vũ Ngǫc [PV09], by the following $5$ symplectic invariants:
@@
be the set of focus-focus critical values, ordered in such a way that $x_{1}\leq x_{2}\leq...\leq x_{m_{f}}$ and let $B_{r}$ be the set of regular values in $B:=F(M)$. For $i\in\{1,\dots,m_{f}\}$ and $\epsilon\in\{-1,+1\}$, define $\mathcal{L}_{i}^{\epsilon}$ to be the vertical ray starting at $c_{i}$ and going to $\pm{\infty}$ depending on the sign of $\epsilon$, i.e., $\mathcal{L}_{i}^{\epsilon}=\{(x_{i},y)\mid\epsilon y\geq\epsilon y_{i}\}$. Given $\vec{\epsilon}=(\epsilon_{1},...,\epsilon_{m_{f}})\in\{-1,+1\}^{m_{f}}$, we define the line segment $l_{i}:=F(M)\cap\mathcal{L}_{i}^{\epsilon_{i}}$ and we set
@@
(Vũ Ngǫc, [Vu07, Theorem 3.8]) Using the notation from above for a semitoric system $(M,\omega,(f_{1},f_{2}))$, for all $\vec{\epsilon}\in\{-1,+1\}^{m_{f}}$, there exists a homeomorphism $f_{\vec{\epsilon}}=(f_{\vec{\epsilon}}^{(1)},f_{\vec{\epsilon}}^{(2)})$ from $B$ to $f_{\vec{\epsilon}}(B)\subseteq{\mathbb{R}}^{2}$ such that
@@
Describing all possible choices (like for instance the signs $\epsilon_{j}$) by means of a group action allows to write the polytope invariant as equivalence class or orbit of this group action. Since we work later always with representatives, we refer for details to Vũ Ngǫc [Vu07].
@@ 2.9. Semitoric systems and almost toric fibrations
Let $(M,\omega,F)$ be a semitoric system. In particular, it induces an almost toric fibration and an almost toric base. Let $\Delta_{\vec{\epsilon}}$ be a representative of the polytope invariant for this semitoric system. Then, by Vũ Ngǫc [Vu07], the eigenline of a nodal value in $\Delta_{\vec{\epsilon}}$ is the vertical line that passes through the node. The affine monodromy can be represented, with respect to an integral affine basis of $\Delta_{\vec{\epsilon}}$, as
@@ 2.10. Symplectic reduction
In this section we recall the following symplectic reduction result from Abreu & Macarini [AM13]: let $x\in\text{int}(P)$ and let $T_{x}:=\mu^{-1}(x)$ denote the corresponding $\mathbb{T}^{N}$ orbit (which is a Lagrangian torus in $M$). Let $K\subset\mathbb{T}^{N}$ be a subtorus of dimension $N-n\geq 0$ determined by an inclusion of Lie algebras $\imath:\mathbb{R}^{N-n}\rightarrow\mathbb{R}^{N}$. The momentum map for the induced action of $K$ on $M$ is given by
@@ 2.11. Displacing toric fibers
In McDuff [McD11] conditions are given to displace Lagrangian toric fibers of toric manifolds. In this section, we summarize the necessary notions and results of McDuff [McD11] which we will need later on.
The affine distance $d_{\text{aff}}(x,y)$ between two points $x$, $y$ on a line $L$ with rational slope is defined as the quotient of the Euclidean distance $d_{E}(x,y)$ and the minimum Euclidean distance from 0 to an integral point $p\in\mathbb{Z}^{n}$ on the line through 0 parallel to $L$. Equivalently, if $\phi$ is any integral affine transformation of $\mathbb{R}^{n}$ that takes $x,y$ to the first coordinate axis (the $x_{1}$-axis) then $d_{\text{aff}}(x,y)=d_{E}(\phi(x),\phi(y))$. An affine line $L=z+\mathbb{R}v$ in $\mathbb{R}^{n}$ is rational if the direction vector $v$ can be taken to be a primitive integral vector in $\mathbb{Z}^{n}$. Given an affine rational line $L$ with primitive direction $v\in\mathbb{Z}^{n}$, the affine distance $d_{\text{aff}}$ between two points $x,y\in L$ is $t\in\mathbb{R}$ such that $x-y=tv\in\mathbb{R}^{n}$. An affine hyperplane $A$ is called rational if it has a primitive integral normal vector $\eta$, i.e., if it is given by an equation of the form $\langle x,\eta\rangle=k$ where $k\in\mathbb{R}$ and $\eta$ is primitive and integral. The affine distance $d_{\lambda}(x,A)$ from a point $x$ to a rational affine hyperplane in the rational direction $\lambda$ is defined as
@@
Let $w$ be a point on a facet $F$, i.e., on a (n1)(n-1)-dimensional face of a rational polytope $\Delta$ and let $\lambda\in\mathbb{Z}^{n}$ be integrally transverse to $F$. The probe $p_{F,\lambda}(w)=:p_{\lambda}(w)$ with direction $\lambda\in\mathbb{Z}^{n}$ and initial point $w\in F$ is the half open line segment consisting of $w$ together with the points in $\text{int}(\Delta)$ that lie on the ray from $w$ in direction $\lambda$.
@@
(McDuff [McD11, Lemma 2.4]) Let $\Delta$ be a smooth momentum polytope associated with a symplectic toric manifold $(M,\omega,F)$, i.e., $\Delta=F(M)$. Let $u\in\text{int}(\Delta)$ lie on the probe $p_{F,\lambda}(w)$. If $w$ lies in the interior of its corresponding facet and if $u$ is less than halfway measured from the boundary along $p_{\lambda}(w)$, then the fiber $F^{-1}(u)$ is displaceable.
@@ 2.12. Special Hamiltonian diffeomorphisms of the disk
In this subsection, we recall a result from Abreu &\& Borman &\& McDuff [ABM14] that concerns Hamiltonian diffeomorphisms of the disk which we will need to displace focus-focus fibers in Section 3.3.
@@
For our purposes, we pick a(s):=Bsa(s):=B-s. For $r>0$, let $\mathbb{D}(r)\subset\mathbb{R}^{2}$ denote the $2$-dimensional disk of radius $r$ centered at the origin.
Lemma 2.24 (Abreu &\& Borman &\& McDuff, [ABM14, Lemma 6.4.2]).
@@
[…]
@@
If $A>\frac{B}{2}$, every circle of radius less than BAB-A is displaced by $\rho$.
@@ 2.13. Symplectic quasi-states
Symplectic quasi-states are a useful tool to find nondisplaceable subsets in a symplectic manifold, and so are the notions of superheavy, heavy and pseudoheavy sets. Entov &\& Polterovich [EP09, EP06] introduced the notions of heaviness and superheaviness. Furthermore, Kawasaki & Orita [KO21] added the notion of pseudoheaviness. The definition of a symplectic quasi-state and pseudoheavy, heavy, superheavy subsets is based on spectral invariants in Hamiltonian Floer theory (see Oh [Oh05]). We now recall in more detail some definitions and results from these papers that we need later on.
@@
  • Quasi-subaditivity: $\zeta(H_{1}+H_{2})\leq\zeta(H_{1})+\zeta(H_{2})$ for all $H_{1},H_{2}\in C^{\infty}(M)$ satisfying $\{H_{1},H_{2}\}=0$, where $\{\cdot,\cdot\}$ is the Poisson bracket induced by the symplectic form.
@@
[…]
@@
Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$. A closed subset $X$ of $M$ is $\zeta\textbf{-heavy}$ (resp. $\zeta\textbf{-superheavy}$) if
@@
for all $H\in C^{\infty}(M)$. A closed subset $X$ is said to be $\zeta\textbf{-pseudoheavy}$ if for all open neighborhood $U$ of $X$ there exists a function $F\in C^{\infty}(U)$ such that $\zeta(F)>0$.
@@
(Kawasaki &\& Orita [KO21, Proposition 1.8], Entov &\& Polterovich [EP09, Theorem $1.4$]) Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$:
@@
(Kawasaki &\& Orita, [KO21, Theorem 1.7]) Let $(M,\omega,F)$ be an integrable system. Let $\zeta:C^{\infty}(M)\rightarrow\mathbb{R}$ be a partial symplectic quasi-state on $(M,\omega)$. Then there exists $y_{0}\in F(M)$ such that $F^{-1}(y_{0})$ is $\zeta$-pseudoheavy.
@@
Entov & Polterovich [EP06] showed the existence of a partial symplectic quasi-state for $2n$-dimensional symplectic manifolds which are rational and strongly semi-positive. Due to the work of Oh [Oh09] and Usher [Ush10, Ush08] the conditions can be dropped, see Entov [Ent14] for more details. Note that all symplectic 4-manifolds are strongly semi-positive.
@@
(Entov &\& Polterovich, [EP06, Theorem 2.1], Usher [Ush10, Ush08], Oh [Oh09]) Any integrable system $(M,\omega,F)$ on a symplectic $4$-manifold $(M,\omega)$ has a nondisplaceable fiber.
@@
(Polterovich &\& Rosen, [PR14, Corollary 6.1.6]) Any stem is nondisplaceable by the group of symplectomorphisms of $(M,\omega)$.
Borman [Bor13] showed that symplectic quasi-states behave well under symplectic reduction on a superheavy subset: let $(W,\omega)$ be a $2n$-dimensional closed symplectic manifold equipped with a smooth map $\Phi=(\Phi_{1},\cdot\cdot\cdot,\Phi_{k}):W\rightarrow\mathbb{R}^{k}$ and a regular level set $Z:=\Phi^{-1}(0)$. Suppose that all $\Phi_{i}$ mutually Poisson commute at each point in $Z$ and that $\Phi$ induces a free Hamiltonian $\mathbb{T}^{k}$-action on $Z$. Let $(M^{\text{red},0}:=Z/\mathbb{T}^{k},\omega^{\text{red},0})$ be the reduced space when performing symplectic reduction at 0. Denote by $\pi:Z\rightarrow M^{\text{red},0}$ the quotient map.
@@
(Borman, [Bor13, Theorem 1.1]) Let $(W,\omega)$ be a $2n$-dimensional closed symplectic manifold and $Z$ has above. If $\zeta:C^{\infty}(W,\omega)\rightarrow\mathbb{R}$ is a symplectic quasi-state satisfying the PB-inequality and $Z$ is superheavy for $\zeta$, then $\zeta$ induces a symplectic quasi-state